OpenStudy (anonymous):

Solve: x^2+5x+12=−2x x^2+7x+12=0

OpenStudy (anonymous):

Use the Order of Operations. PEMDAS.

OpenStudy (anonymous):

1.Parenthesis 2. Exponents 3. Multiply 4. Divide 5. Add 6. Subtract

OpenStudy (solomonzelman):

x^2+5x+12=−2x you added 2x to both sides, and got x^2+7x+12=0 then `x^2+7x+12` = `(x+4)(x+3)`

OpenStudy (anonymous):

So is that the answer? @SolomonZelman

OpenStudy (solomonzelman):

I gave you enough help to solve the problem, don't just ask me for the answer !!!